OpenStudy (anonymous):

(2y^3 + 3y^2 - 4y - 5) / (y + 1)

OpenStudy (aroub):

(2y^3 + 3y^2 - 4y - 5) ------------------ (y+1) = (y+1) (2y^2+y-5) -------------- (y+1) = 2y^2+y-5